Solve the recurrence relation an=6 an-2-12 an-2 + 8 an-2 + n2 2n
1
Expert's answer
2021-09-21T02:12:13-0400

Let us solve the recurrence relation "a_n=6 a_{n-2}-12a_{n-2} + 8a_{n-2} + n^2+ 2n," which is equivalent to "a_n-2a_{n-2}= n^2+ 2n." The characteristic equation "k^2-2=0" of the homogeneous recurence relation "a_n-2a_{n-2}=0" has the roots "k_1=\sqrt{2},\ k_2=-\sqrt{2}."

It follows that the solution of "a_n-2a_{n-2}= n^2+ 2n" is of the form "a_n=C_1(\sqrt{2})^n+C_2(-\sqrt{2})^n+b_p(n)," where "b_p(n)=an^2+bn+c." Then "an^2+bn+c-2(a(n-2)^2+b(n-2)+c)=n^2+2n," and hence "an^2+bn+c-2(an^2-4an+4a+bn-2b+c)=n^2+2n." It follows that "-an^2+(8a-b)n-8a+4b-c=n^2+2n," and thus "a=-1,\ 8a-b=2,\ -8a+4b-c=0." Therefore, "a=-1,\ b=-10,\ c= -32."


We conclude that the solution is the following:

"a_n=C_1(\sqrt{2})^n+C_2(-\sqrt{2})^n-n^2-10n-32."